40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 25 02:03:59 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts 
    Reply  Fri Jul 28 15:36:32 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts errSig.pdfQPD_simulink.pngfeedbackTopology.pdf
       Reply  Tue Aug 1 16:05:01 2017, gautam, Update, Optical Levers, Optical lever tuning - cost function construction loopPerformance.pdfOplevLoop.pngOL_errSigs.pdfDAC_saturation.pdf
          Reply  Thu Nov 16 13:57:01 2017, gautam, Update, Optical Levers, Optical lever noise VertexOLnoise.pdfITMYvsSRM.pdf
             Reply  Thu Nov 16 15:43:01 2017, rana, Update, Optical Levers, Optical lever noise 
                Reply  Tue Nov 21 11:37:29 2017, gautam, Update, Optical Levers, BS OL calibration updated OL_calib_BS_PERROR.pdfOL_calib_BS_YERROR.pdfVertexOLnoise_updated.pdf
                   Reply  Tue Nov 21 16:28:23 2017, gautam, Update, Optical Levers, BS OL calibration updated BS_oplev_sweep.png
                      Reply  Tue Nov 21 23:04:12 2017, gautam, Update, Optical Levers, Oplev "noise budget" vertexOL_noises.pdf
                         Reply  Wed Nov 22 05:41:32 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                            Reply  Wed Nov 22 15:29:23 2017, gautam, Update, Optical Levers, Oplev "noise budget" OplevTest.jpg
                               Reply  Wed Nov 22 16:40:00 2017, Koji, Update, Optical Levers, Oplev "noise budget" 
                               Reply  Wed Nov 22 19:20:01 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                                  Reply  Wed Nov 22 23:56:14 2017, gautam, Update, Optical Levers, Oplev "noise budget" 
                                     Reply  Thu Nov 23 18:03:52 2017, gautam, Update, Optical Levers, Oplev "noise budget" ITMY_P_noise.pdfITMY_Y_noise.pdf
Message ID: 13452     Entry time: Wed Nov 22 23:56:14 2017     In reply to: 13451     Reply to this: 13453
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: Oplev "noise budget" 

Do not turn on BS/ITMY/SRM/PRM Oplev servos without reading this elog and correcting the needful!

I've setup a test setup on the ITMY Oplev table. Details + pics to follow, but for now, be aware that

  1. I've turned off the HeNe that is used for the SRM and ITMY Oplev.
  2. Moved one of the HeNe's Steve setup on the SP table to the ITMY Oplev table.
  3. Output power was 2.5mW, whereas normal power incident on this PD was ~250uW.
  4. So I changed all transimpedance gains on the ITMY Oplev QPD from 100k to 10k thin film - these should be changed back when we want to use this QPD for Oplev purposes again. Note that I did not change the compensation capacitors C3-C6, as with 10k transimpedance, and assuming they are 2.2nF, we get a corner frequency of 6.7kHz. The original schematic recommends 0.1uF. In hindsight, I should have changed these to 22nF to keep roughly the same corner frequency of ~700Hz.
    I've implemented this change as of ~5pm Nov 23 2017 - C3-C6 are now 22nF, so the corner frequency is 676Hz, as opposed to 723Hz before... This should also be undone when we use this QPD as an Oplev QPD again...
  5. I marked the position of the ITMY Oplev QPD with sharpie and also took pics so it should be easy enough to restore when we are done with this test.
  6. I couldn't get the HeNe to turn on with any of the power supplies I found in the cabinet, so I borrowed the one used to power the BS/PRM. So these oplevs are out of commission until this test is done.
  7. There is a single steering mirror in a Polaris mount which I used to center the spot on the QPD.
  8. The specular reflection (~250uW, i.e. 10% of the power incident on the QPD) is dumped onto a clean razor beam stack. Steve can put in a glass beam dump on Monday.
  9. Just in case someone accidentally turns on some servos - I've disabled the inputs to the BS, PRM and SRM oplevs, and set the limiter on the ITMY servo to 0.

Here are some pics of the setup: https://photos.app.goo.gl/DHMINAV7aVgayYcf1.None of the existing Oplev input/output steering optics were touched. Steve can make modifications as necessary, perhaps we can make similar mods to the SRM Oplev QPD and the BS one to run the HeNe test for a few days...

Quote:

too complex; just shoot straight from the HeNe to the QPD. We lower the gain of the QPD by changing the resistors; there's no sane reason to keep the existing 100k resistors for a 2 mW beam. The specular reflection of the QPD must be dumped on a black glass V dump (not some flimsy anodized aluminum or dirty razor stack)

 

ELOG V3.1.3-